# SonicSim: A customizable simulation platform for speech processing in moving sound source scenarios

### v1

@@ Abstract
The systematic evaluation of speech separation and enhancement models under moving sound source conditions typically requires extensive data comprising diverse scenarios. However, real-world datasets often contain insufficient data to meet models’ training and evaluation requirements. Although synthetic datasets offer a larger volume of data, their acoustic simulations lack realism. Consequently, neither real-world nor synthetic datasets effectively fulfill practical needs. To address these issues, we introduce SonicSim, a synthetic toolkit designed to generate highly customizable data for moving sound sources. SonicSim is developed based on the embodied AI simulation platform, Habitat-sim, supporting multi-level adjustments, including scene-level, microphone-level, and source-level, thereby generating more diverse synthetic data. Leveraging SonicSim, we constructed a moving sound source benchmark dataset, SonicSet, using LibriSpeech, Freesound Dataset 50k (FSD50K) and Free Music Archive (FMA), and 90 scenes from the Matterport3D to evaluate speech separation and enhancement models. Additionally, to investigate the differences between synthetic and real-world data, we randomly selected 5 hours of raw data without reverberation from the SonicSet validation set to record a real-world speech separation dataset, which set a reference for comparing SonicSet and other synthetic datasets. Similarly, we utilized the real-world speech enhancement dataset RealMAN to validate the acoustic gap between the SonicSet dataset and other synthetic datasets for speech enhancement. The results indicate that models trained on SonicSet generalized better to real-world scenarios compared with other synthetic datasets. Demo and code are publicly available at https://cslikai.cn/SonicSim/.
@@ 1 Introduction
Speech separation (Cherry, 1953) and enhancement (Loizou, 2007) are classic tasks in speech signal processing, with the goal of extracting the target speech signal from noisy audio. In recent years, the release of large-scale synthetic datasets, such as LibriMix (Cosentino et al., 2020) and the DNS Challenge (Reddy et al., 2020), has contributed to the improvement of speech separation and enhancement models. Many methods Wang et al. (2023); Subakan et al. (2021); Chen et al. (2020a); Luo & Yu (2023b); Li et al. (2022e); Li & Luo (2023) have demonstrated the ability to transfer across different test environments in zero-shot settings. Applications like online meetings (Rao et al., 2021) and human-computer interaction (Yu & Chen, 2017) have greatly benefited from advancements in these state-of-the-art (SOTA) speech separation and enhancement methods. However, despite the remarkable progress, there remains a significant acoustic mismatch between synthetic data and real-world data (Yang et al., 2024). This discrepancy often results in suboptimal model performance in real-world scenarios when trained on synthetic data (Sivaraman & Kim, 2022).
@@
In previous studies, efforts to reduce the gap between synthetic and real-world data have primarily focused on more accurately modeling room reverberation. Current datasets, such as WHAMR! (Maciejewski et al., 2020) and the DNS Challenge (Reddy et al., 2020), simulate room impulse responses (RIRs) using methods like finite element analysis (Jarrett et al., 2012), ray tracing (Lopez-Hernandez et al., 2000), or the image source method (Luo & Yu, 2023a). However, RIRs generated by these methods still exhibit significant acoustic discrepancies when compared to real-world reverberation. Specifically, from the perspective of RIR simulation, existing approaches face three major challenges: 1) Occlusion of obstacles. RIRs generated based on the image source method struggle to effectively handle varying quantities and shapes of obstacles within a room (Aralikatti et al., 2022). This issue becomes particularly pronounced when objects obstruct the sound path between the source and the microphone, making it challenging to accurately reproduce their acoustic effects. 2) Complex room geometry. Simulated RIRs typically assume that rooms have cuboid-shaped geometries (Southern et al., 2011), which is a significant deviation from real-world complex environments, such as lecture halls with tiered seating. This discrepancy leads to less realistic simulation outcomes. 3) Complicated room surfaces and object materials. Different materials have distinct sound absorption, reflection, and scattering properties, and when multiple materials are present in a scene, accurately replicating such complex acoustic environments in simulations becomes difficult. For example, hard walls generally reflect more sound, while soft materials like carpets or furniture absorb more. Therefore, an accurate simulation of RIRs remains a challenging problem.
@@
Additionally, a notable characteristic of these datasets is that they are primarily collected from static sound sources (Cosentino et al., 2020; Maciejewski et al., 2020; Reddy et al., 2020), which makes these datasets more suitable for non-moving scenarios such as online meetings. In moving sound source scenarios, speakers often move around within a space while static microphones capture their speech, a setup more relevant to applications such as robotic navigation (Wang et al., 2004). Compared to the collection of static speech datasets, capturing large-scale moving sound source datasets is more time-consuming (Yang et al., 2024). As a result, there is a relative scarcity of data on moving sound source scenarios, which limits the development and evaluation of speech separation and enhancement models in such tasks. Even if the cost of data acquisition can be justified and the data is sufficiently accurate, real-world datasets are still constrained by the availability of physical environments. For example, for specific indoor layouts, it is necessary to collect data directly from real-world scenes. Once collected, the distribution of real-world datasets is often fixed, making them difficult to adjust flexibly. This limits researchers’ ability to conduct personalized experiments and increases the risk of models overfitting to a specific dataset, thereby rendering benchmarks outdated. Thus, despite the high research value of moving sound source scenarios, the lack of relevant datasets significantly hinders deeper exploration in this area.
@@
On the SonicSet dataset, we conducted extensive quantitative experiments on 11 speech separation methods and 9 speech enhancement methods, thoroughly analyzing the performance of each approach. To evaluate the gap between SonicSet and real-world environments, we randomly selected some raw audio without reverberation from the SonicSet validation set and recorded it in real-life scenarios, constructing a speech separation dataset consisting of 10 scenes with a total duration of 5 hours. Additionally, for the speech enhancement task, we utilized the RealMAN test set (Yang et al., 2024), which includes recordings of moving sound sources from real-world environments, to evaluate the gap between the synthetic dataset and real environments. The experimental results demonstrated that models trained on the SonicSet dataset generalized to real-world environments, validating the effectiveness and potential of SonicSim for speech research.
@@ 2 Related works
Synthetic datasets: Synthetic datasets (Reddy et al., 2020; Cosentino et al., 2020; Maciejewski et al., 2020; Chen et al., 2020b) provide an alternative solution for speech separation and enhancement tasks, typically generating synthetic data by convolving RIR with multiple source signals or noise signals to ensure realism. These simulated RIRs are often generated using the image source methods Allen & Berkley (1979). Datasets such as WHAMR! Maciejewski et al. (2020) introduce synthetic reverberation and noise environments, increasing the complexity of the data. However, synthetic datasets are often created under tightly controlled simulation conditions, which reduces the authenticity of the data and cannot reflect the performance of speech separation and enhancement methods in real-world environments. In contrast, SonicSet not only ensures the acoustic plausibility of simulated audio (e.g., occlusions and material properties) but also supports a broader range of customizable options (e.g., varying geometric structures and multiple microphone array configurations). This provides more challenging and realistic environments for speech separation and enhancement methods.
Reverberation simulation tools: In recent years, significant progress has been made in the development of reverberation simulation tools. FRAM-RIR (Luo & Yu, 2023a) offers a rapid RIR simulation method using a stochastic approximation of the image source method, which improves simulation speed. RIR-Generator (Tang et al., 2020) is an open-source tool based on the image source method, allowing users to set parameters. Pyroomacoustics (Scheibler et al., 2018) provides a flexible Python library that enables fast simulation of source-microphone setups in 2D or 3D rooms. For more complex acoustic environments, COMSOL Multiphysics (Multiphysics, 1998) combines finite element methods to provide detailed modeling of a room’s acoustic characteristics. Compared to these tools, SonicSim based on Habitat-sim (Savva et al., 2019) surpasses existing simulation tools in acoustic realism, making synthetic audio closely resembling recordings in real-world environments. Additionally, SonicSim provides a range of practical features that allow users to efficiently create diverse acoustic scenes tailored to specific needs—capabilities that most existing reverberation simulation tools lack.
@@ 3.1 Customizable data generator: SonicSim
The customizable dataset generator, SonicSim, is specifically designed to generate datasets for speech tasks. Built on Habitat-sim (Savva et al., 2019), SonicSim leverages its highly realistic audio renderer (demonstrated in (Chen et al., 2022a)) and high-performance 3D simulator to produce high-quality audio data adapted to various acoustic environments. SonicSim provides a rich set of annotations, including source and microphone position maps, clean audio, and audio with reverberation and noise, without incurring additional data collection costs. More importantly, SonicSim provides users with extensive control over the dataset generation process, allowing customization of scene layouts, scene materials, source and microphone positions, and microphone types, while ensuring physical realism through its physics engine. By adjusting these configurations, SonicSim can flexibly generate diverse acoustic environment data to meet the specific requirements of various tasks. The following main functions are included in SonicSim, as shown in Figure 1.
@@
3D scene import: Through Habitat-sim (Savva et al., 2019), SonicSim can import various realistic 3D assets generated through simulations or scans, such as those from the Matterport3D dataset (Chang et al., 2017). SonicSim leverages Habitat-sim111https://aihabitat.org/docs/habitat-sim/habitat_sim.sim.SimulatorConfiguration.html to interpret metadata and structural information from external datasets and convert them into Habitat-sim’s native scene format. Geometric data, material properties, and semantic annotations are transformed to ensure that the imported scenes maintain their original high fidelity. This ensures consistency within the simulation environment. This workflow not only simplifies the process but also makes the generation of complex and realistic acoustic environments more efficient and scalable, enabling researchers to conduct experiments in environments that closely resemble real-world scenarios.
@@
Acoustic environment simulation: SonicSim uses Habitat-sim to simulate various acoustic features in 3D environments222https://aihabitat.org/docs/habitat-sim/habitat_sim.sensor.AudioSensor.html. Its capabilities include: 1) accurately simulating sound reflections within room geometries using indoor acoustic modeling and bidirectional path tracing algorithms; 2) mapping the semantic labels of 3D scenes to material properties to set the acoustic features of different surfaces, such as absorption, scattering, and transmission coefficients; 3) extended Habitat-sim acoustic simulation capabilities enable the synthesis of moving sound source data based on sound source paths. SonicSim can generate audio that closely resembles real-world conditions, ensuring the physical realism of acoustic simulations.
@@
Microphone types: Habitat-sim offers comprehensive microphone configurations, supporting various audio formats such as mono, binaural, and ambisonics333https://github.com/facebookresearch/habitat-sim/blob/main/docs/AUDIO.md. Additionally, we have integrated common linear and circular microphone arrays, allowing users to customize the shape of microphone arrays to meet different experimental requirements. Specifically, this functionality is realized by arranging multiple mono audio sensors within a 3D space. We provide a flexible API that enables researchers to represent the desired array layout by inputting custom functions or data structures. The simulation engine then instantiates microphones in the environment based on the input, ensuring precise acoustic modeling of the array geometry.
@@ 3.2 Moving sound source dataset: SonicSet
We used SonicSim to construct a moving sound source dataset, SonicSet, specifically designed for research on moving speech separation and enhancement tasks. This dataset supports diverse acoustic scenarios by simulating microphones, sound sources, and noise sources randomly placed in the environment. A detailed data analysis can be found in Appendix A. As shown in Figure 3, the entire data generation process is outlined, ensuring the validity of the sound source and microphone configurations under varying positions and dynamic changes.
@@ 3.2.1 Data source
For the audio data, we primarily used speech data from the LibriSpeech dataset (Panayotov et al., 2015), and noise data from the FSD50K (Fonseca et al., 2021) and the FMA datasets (Defferrard et al., 2017). LibriSpeech is an English speech dataset consisting of approximately 1,000 hours of 16 kHz sampled read speech, sourced from audiobooks in the LibriVox project. We selected train-clean-360 as the training set for speech content, dev-clean as the validation set, and test-clean for testing. FSD50K is an open dataset of manually labeled sound events, containing 51,197 audio clips with a total duration of over 100 hours, covering 200 sound categories. We downsampled them to 16 kHz using the Librosa (McFee et al., 2015). The dataset was divided into training, validation, and test sets in a 7:2:1 ratio. For the FMA dataset, an open music dataset designed to support a variety of tasks in the music domain, we employed a pre-trained BSRNN model (Luo & Yu, 2023b) on music separation tasks to remove vocals from the tracks, ensuring that the speech separation process would not be affected during data synthesis. Subsequently, we downsampled the music data to 16 kHz using Librosa. Similar to FSD50K, the FMA dataset was divided into training, validation, and test sets following a 7:2:1 ratio.
@@ 3.2.2 Dataset construction
The audio simulation pipeline in the SonicSet dataset, as illustrated in Figure 3, could be divided into the following steps. First, we selected a 3D scene from the Matterport3D dataset and imported it into SonicSim to initialize the acoustic environment. Second, we randomly selected the placement positions for the microphone and the sound source from a certain layer of the scene. Once the microphone’s position was determined, the sound source’s starting position and the noise source’s position were randomly chosen within a 1-8 meter radius from the microphone. Third, based on the initial position of the sound source, we selected a sound source’s endpoint within a 1-8 meter range from both the microphone and the sound source’s initial position. We then employed SonicSim’s trajectory function to calculate and generate a movement path for the sound source. This trajectory function calculated the distance between consecutive points and converts them into a sampling rate, dividing the total number of samples for the dynamic audio based on these distances, and subsequently calculating interpolation indices and weights. Fourth, SonicSim calculated the corresponding RIRs for the different positions along the path and convolved the source audio with these RIRs. Finally, based on the precomputed indices and weights, we extracted audio segments corresponding to the start and end positions of each section from the convolution output and mixed them according to the interpolation weights. Through this process, a temporally coherent audio signal was generated, accurately reflecting the movement of the sound source within the spatial environment.
@@
To create realistic synthetic mixed audio, we set the length of each mixed audio clip to 60 seconds. Each speech segment was composed of 3-5 full audio clips with the same speaker randomly selected from LibriSpeech, arranged into a 60-second audio sequence by randomly selecting start positions within 0-8 seconds for each clip. This approach ensured various overlap rates in the mixed audio while utilizing SonicSim for moving sound source synthesis. For noise data, 6-8 segments of environmental noise were randomly selected from the divided FSD50K dataset, and 6-8 segments of musical noise were randomly chosen from the divided FMA dataset. These noise segments were arranged into 60-second audio clips by choosing random start positions within 0-4 seconds. We used Loudness Units relative to Full Scale (LUFS) to adjust volume levels, setting the speech at -17 LUFS, environmental noise at -21 LUFS, and musical noise at -24 LUFS. When generating the mixed audio, different speakers were randomly chosen from the speech data, and a segment of environmental noise was randomly selected to create a mixed audio clip with background noise, while a segment of musical noise was randomly chosen to create a mixed audio clip with a musical noise. These mixed audio samples were used to test the robustness of speech separation and enhancement models in various noisy environments. For each dataset group (three speech clips + one environmental noise clip + one musical noise clip), we simulated within the same scene, where the speech clips were simulated as moving sources and the noise was simulated as static sources. We present an example of SonicSet in Figure 5 in Appendix B and provide the corresponding audio metadata in the JSON file 1 in Appendix B.

### v2

@@ Abstract
Systematic evaluation of speech separation and enhancement models under moving sound source conditions requires extensive and diverse data. However, real-world datasets often lack sufficient data for training and evaluation, and synthetic datasets, while larger, lack acoustic realism. Consequently, neither effectively meets practical needs. To address this issue, we introduce SonicSim, a synthetic toolkit based on the embodied AI simulation platform Habitat-sim, designed to generate highly customizable data for moving sound sources. SonicSim supports multi-level adjustments—including scene-level, microphone-level, and source-level—enabling the creation of more diverse synthetic data. Leveraging SonicSim, we constructed a benchmark dataset called SonicSet, utilizing LibriSpeech, Freesound Dataset 50k (FSD50K), Free Music Archive (FMA), and 90 scenes from Matterport3D to evaluate speech separation and enhancement models. Additionally, to investigate the differences between synthetic and real-world data, we selected 5 hours of raw, non-reverberant data from the SonicSet validation set and recorded a real-world speech separation dataset, providing a reference for comparing SonicSet with other synthetic datasets. For speech enhancement, we utilized the real-world dataset RealMAN to validate the acoustic gap between SonicSet and existing synthetic datasets. The results indicate that models trained on SonicSet generalize better to real-world scenarios compared to other synthetic datasets. Code is publicly available at https://cslikai.cn/SonicSim/.
@@ 1 Introduction
Speech separation (Cherry, 1953) and enhancement (Loizou, 2007) are classic tasks focused on extracting target speech from noisy audio. The release of large-scale synthetic datasets like LibriMix (Cosentino et al., 2020) and the DNS Challenge (Reddy et al., 2020) has propelled model advancements, with many methods (Chen et al., 2020a; Luo & Yu, 2023b) demonstrating cross-environment transferability in zero-shot settings. These SOTA methods have significantly benefited applications such as online meetings (Rao et al., 2021) and human-computer interaction (Yu & Chen, 2017). However, a significant acoustic mismatch between synthetic and real-world data persists, leading to suboptimal model performance in real-world scenarios (Sivaraman & Kim, 2022).
@@
Previous efforts to bridge the gap between synthetic and real-world data have mainly focused on more accurate modeling of room reverberations. Datasets like WHAMR! (Maciejewski et al., 2020) and the DNS Challenge (Reddy et al., 2020) simulate room impulse responses (RIRs) using methods such as finite element analysis (Jarrett et al., 2012), ray tracing (Lopez-Hernandez et al., 2000), or the image source method (Luo & Yu, 2023a). However, from the perspective of RIR simulation, existing approaches face three major challenges: 1) Occlusion of obstacles: RIRs based on the image source method struggle with varying quantities and shapes of obstacles, particularly when objects obstruct the sound path between the source and microphone (Aralikatti et al., 2022). 2) Complex room geometry: Simulations often assume cuboid-shaped rooms (Southern et al., 2011), deviating from complex real-world environments, such as lecture halls with tiered seating. 3) Complicated room surfaces and object materials: Different materials have distinct acoustic properties, and accurately replicating scenes using multiple materials is challenging. Therefore, accurate simulation of RIRs remains a difficult problem.
@@
Existing datasets are primarily collected from static sound sources (Cosentino et al., 2020; Maciejewski et al., 2020; Reddy et al., 2020), making them suitable for non-moving scenarios like online meetings. In contrast, moving sound source scenarios—where speakers move around while static microphones capture their speech—are more relevant to applications such as robotic navigation (Wang et al., 2004). However, capturing large-scale datasets for moving sound sources is more time-consuming, leading to a scarcity of such data, which limits the development and evaluation of speech separation and enhancement models in these tasks. Additionally, real-world datasets are constrained by physical environments and, once collected, their distributions are fixed, making them difficult to adjust flexibly and increasing the risk of models overfitting to specific datasets. Thus, despite their high research value, the lack of relevant datasets for moving sound source scenarios significantly hinders deeper exploration in this area.
@@
On the SonicSet dataset, we conducted extensive quantitative experiments on 11 speech separation methods and 9 speech enhancement methods, thoroughly analyzing the performance of each approach. To assess the gap between SonicSet and real-world environments, we created a speech separation dataset by recording raw, non-reverberant audio from the SonicSet validation set in real-life scenarios across 10 scenes, totaling 5 hours. For the speech enhancement task, we utilized the RealMAN test set (Yang et al., 2024), which includes real-world recordings of moving sound sources, to evaluate discrepancies between the synthetic dataset and real environments. The experimental results demonstrated that models trained on SonicSet generalized well to real-world conditions, confirming the effectiveness and potential of SonicSim for speech research.
@@ 2 Related works
Synthetic datasets: Synthetic datasets (Reddy et al., 2020; Cosentino et al., 2020; Maciejewski et al., 2020; Chen et al., 2020b) generate data by convolving Room Impulse Responses (RIRs) with multiple source or noise signals to enhance realism, often using image source methods (Allen & Berkley, 1979). While datasets like WHAMR! (Maciejewski et al., 2020) introduce synthetic reverberation and noise, they are created under controlled conditions, limiting their authenticity and inability to fully reflect real-world performance of speech separation and enhancement methods. In contrast, SonicSet ensures acoustic plausibility and offers a wider range of customizable options, such as varying geometric structures and multiple microphone array configurations, providing more challenging and realistic environments for these tasks.
Reverberation simulation tools: Recent advances in reverberation simulation tools have significantly contributed to the field. FRAM-RIR (Luo & Yu, 2023a) uses a stochastic approximation of the image source method for rapid RIR simulation. RIR-Generator (Tang et al., 2020) allows parameter settings based on the image source method, and Pyroomacoustics (Scheibler et al., 2018) offers a flexible Python library for fast simulation in 2D or 3D rooms. For complex acoustic environments, COMSOL Multiphysics (Multiphysics, 1998) employs finite element methods for detailed room modeling. Compared to these tools, SonicSim based on Habitat-sim (Savva et al., 2019) surpasses others in acoustic realism, producing synthetic audio that closely resembles real-world recordings. It also provides practical features for efficiently creating diverse acoustic scenes tailored to specific needs—capabilities lacking in most existing reverberation simulation tools.
@@ 3.1 Customizable data generator: SonicSim
The customizable dataset generator, SonicSim, is specifically designed to generate datasets for speech tasks. Built on Habitat-sim (Savva et al., 2019), SonicSim leverages its highly realistic audio renderer (demonstrated in (Chen et al., 2022a)) and high-performance 3D simulator to produce high-quality audio data adapted to various acoustic environments. SonicSim provides a rich set of annotations, including source and microphone position maps, clean audio, and audio with reverberation and noise, without incurring additional data collection costs. More importantly, SonicSim provides users with extensive control over the dataset generation process, allowing customization of scene layouts, scene materials, source and microphone positions, and microphone types, while ensuring physical realism through its physics engine. By adjusting these configurations, SonicSim can flexibly generate diverse acoustic environment data to meet the specific requirements of various tasks. In Appendix A, we discussed in detail the impact of environmental factors on model performance. The following main functions are included in SonicSim, as shown in Figure 1.
@@
3D scene import: Through Habitat-sim (Savva et al., 2019), SonicSim can import various realistic 3D assets generated through simulations or scans, such as those from the Matterport3D dataset (Chang et al., 2017). SonicSim leverages Habitat-sim111https://aihabitat.org/docs/habitat-sim/habitat_sim.sim.SimulatorConfiguration.html to interpret metadata and structural information from external datasets and convert them into Habitat-sim’s native scene format. Geometric data, material properties, and semantic annotations are transformed to ensure that the imported scenes maintain their original high fidelity. Importing 3D scenes with various shapes, such as cylinders, enables Habitat-sim to simulate different acoustic properties according to different spatial layouts. SonicSim allows users to add additional objects to the scene; however, the 3D scenes themselves cannot be modified by users, as they are fixed during the import process.
@@
Acoustic environment simulation: SonicSim uses Habitat-sim to simulate various acoustic features in 3D environments222https://aihabitat.org/docs/habitat-sim/habitat_sim.sensor.AudioSensor.html. Its capabilities include: 1) accurately simulating sound reflections within room geometries using indoor acoustic modeling and bidirectional path tracing algorithms. This effectively simulates the acoustic effect of different objects blocking sound.; 2) mapping the semantic labels of 3D scenes to material properties to set the acoustic features of different surfaces, such as absorption, scattering, and transmission coefficients. SonicSim supports a wide range of materials and objects, and users can modify the physical and categorical attributes of materials; 3) extended Habitat-sim acoustic simulation capabilities enable the synthesis of moving sound source data based on sound source paths.
@@
Microphone types: Habitat-sim offers comprehensive microphone configurations, supporting various audio formats such as mono, binaural, and ambisonics333https://github.com/facebookresearch/habitat-sim/blob/main/docs/AUDIO.md. Additionally, we have integrated common linear and circular microphone arrays, allowing users to customize the shape of microphone arrays to meet different experimental requirements. Specifically, we first define multiple single-channel receivers corresponding to the number of microphones in the array. Next, we configure the spatial relative coordinates of these single-channel receivers based on the desired array shape. Finally, we bind these receivers together and adjust the absolute coordinates to set the position of the microphone array within the environment. We provided a flexible API that enables researchers to represent the desired array layout by inputting custom functions.
@@ 3.2 Moving sound source dataset: SonicSet
We used SonicSim to construct a moving sound source dataset, SonicSet, specifically designed for research on moving speech separation and enhancement tasks. This dataset supports diverse acoustic scenarios by simulating microphones, sound sources, and noise sources randomly placed in the environment. A detailed data analysis can be found in Appendix B. As shown in Figure 3, the entire data generation process is outlined, ensuring the validity of the sound source and microphone configurations under varying positions and dynamic changes.
@@ 3.2.1 Data source
For the audio data, we used speech from the LibriSpeech dataset (Panayotov et al., 2015) and noise from FSD50K (Fonseca et al., 2021) and FMA (Defferrard et al., 2017). LibriSpeech consists of approximately 1,000 hours of 16 kHz read speech from audiobooks; we selected train-clean-360 for training, dev-clean for validation, and test-clean for testing. FSD50K, containing 51,197 manually labeled sound events totaling over 100 hours and covering 200 categories, was downsampled to 16 kHz using Librosa (McFee et al., 2015) and split into training, validation, and test sets in a 7:2:1 ratio. For FMA, an open music dataset, we used a pre-trained BSRNN model (Luo & Yu, 2023b) to remove vocals, ensuring speech separation was unaffected during data synthesis. We then downsampled the music to 16 kHz and divided it similarly into training, validation, and test sets.
@@ 3.2.2 Dataset construction
The audio simulation pipeline in the SonicSet dataset, illustrated in Figure 3, consists of the following steps. First, we selected a 3D scene from the Matterport3D dataset and imported it into SonicSim to initialize the acoustic environment. Second, we randomly chose positions for the microphone and the sound source within a certain layer, with the sound source’s starting position and the noise source’s position within a 1–8 meter radius from the microphone. Third, based on the sound source’s initial position, we selected an endpoint within a 1–8 meter range from both the microphone and the initial position, using SonicSim’s trajectory function444The trajectory is generated using Habitat-sim’s API, specifically the “habitat_sim.ShortestPath” method. to generate a movement path. Fourth, SonicSim computed the RIRs along the trajectory and convolved them with the source audio. Let denote the source audio signal of duration , and let represent the RIRs at discrete positions , where indexes the audio channels. The convolution at each position is . To achieve realistic spatial movement, we interpolated the convolution results of adjacent positions using a time-varying interpolation weight , calculated based on the Euclidean distances between the current position of the moving source and the adjacent positions: where denotes the Euclidean distance between positions and . This ensures transitions smoothly from to as the source moves from to . The interpolated audio signal is then computed as dynamic acoustic changes as the source moves (Slaney et al., 1996; Freeland et al., 2002). Finally, interpolated audio segments were concatenated to form a moving source. SonicSim adjusts the source’s movement speed based on trajectory length to ensure compatibility with the fixed duration .
@@
To create realistic synthetic mixed audio, we set each mixed audio clip to 60 seconds. Speech segments consisted of 3-5 full audio clips from the same LibriSpeech speaker, arranged into a 60-second sequence with random start positions within 0-8 seconds to ensure varied overlap rates, utilizing SonicSim for moving sound source synthesis. For noise data, we randomly selected 6-8 environmental noise segments from FSD50K and 6-8 musical noise segments from FMA, arranging them into 60-second clips with random start positions within 0-4 seconds. Volume levels were adjusted using Loudness Units relative to Full Scale (LUFS): speech at -17 LUFS, environmental noise at -21 LUFS, and musical noise at -24 LUFS. Mixed audio was generated by randomly choosing different speakers and adding environmental or musical noise, with background and musical noise samples including diverse, time-varying content. In each dataset group (three speech clips plus one environmental noise clip and one musical noise clip), simulations were conducted within the same scene, simulating speech as moving sources and noise as static sources. An example of SonicSet is presented in Figure 5, with corresponding audio metadata and generation hyperparameters provided in JSON files 1 and 2 in Appendix C.